OpenStudy (anonymous):

Integrate \[\int\limits \frac{e ^{2x}}{e ^{x}+3}dx\]

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

this is better let e^x+3=t

OpenStudy (anonymous):

how can be the answer is \[e ^{x}+3\ln \left|e^{x}+3 \right|+C\]???

OpenStudy (anonymous):

Numerator = (t-3)dt Denominator= t

OpenStudy (anonymous):

this method is Integration by substitution Have U studied about it before?

OpenStudy (anonymous):

yup. Let => \[u = e ^{x}+3\] \[du = e ^{x}\] \[=\int\limits \frac{1}{u}\] \[= \ln u + C\] \[= \ln \left| e ^{x}+3 \right|+C\]

OpenStudy (anonymous):

u droped one term!

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

i know i have a mistake

OpenStudy (anonymous):

i dont know how to correct it.

OpenStudy (anonymous):

well \[\frac{e^{2x}dx}{e^x+3}=\frac{e^{x}e^{x}dx}{e^x+3}\\ t=e^x+3 \\ dt=e^x dx\] then u have \[\frac{(t-3)dt}{t}\]

OpenStudy (anonymous):

now try again

OpenStudy (anonymous):

i can't get it. :( please show me.

OpenStudy (anonymous):

is that ok till there?

OpenStudy (anonymous):

are you typing the solution @mukushla ???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\int\limits \frac{e^{2x}}{e^x+3}dx \\ \] then we use a substitution \[t=e^x+3 \\dt= e^x dx \\ e^x=t-3\] now we have \[\int\limits \frac{e^{2x}}{e^x+3}dx= \int\limits \frac{e^{x}e^{x}dx}{e^x+3}=\int\limits \frac{(t-3)}{t}dt=\int\limits (1-\frac{3}{t})dt=t-3 \ ln \left| t \right|+C\]

OpenStudy (anonymous):

Thanks a lot @mukushla .

OpenStudy (anonymous):

ur welcome